LPC meeting summary 02-05-2016 - final
Main purpose of the meeting: Discuss the ongoing intensity ramp up and the programme for the VdM scan.
In the introduction Jamie summarised the status of the commissioning. Due to the problems with the SPS dump the length of the trains in the SPS is currently limited to 72b. This would allow to fill the machine with up to 2000 bunches but would require 30 injections.
Currently the crossing angle of IP1/5 is not correctly diplayed and logged. This is related to the implemntation of the Totem Bump and will need some time to be fixed.
A new service on the lpc website has been put in place. It allows to retrieve the filling schemes for a given fill which reached Stable Beams. Documentation is available on the LPC web-site.
In view of the problems of the week (weasel induced electrical problems and SPS dump problem) a possible re-scheduling of the VdM scan was briefly discussed:
Witold gave an overview on the VdM programme foreseen for the scan in Week 20. Ideally, the programme could be absolved in 2 shifts with two fills. Witold clarified that before the VdM scan there should be a physics fill in the machine so that the luminometers are correctly conditioned. (Loss maps would have to be performed ahead of time.)
There is also some preparation from the machine side still pending. It was mentioned that INDIVs with an emittance of 3 - 3.5µm would have to be prepared in the injectors. Also the DOROS BPMs would have to be checked since they are important for the scan.
LHCb will measure the p-He cross section with the SMOG system when scans are performed in other IPs.
During the first part of the CMS scan, ATLAS will stay head-on since this time will be used to take calibration data for the detectors with individual bunches (the calibration differs between trains and INDIVs.). However during the second part of the CMS scan, where CMS takes beam imaging data and checks the reproducibility, ATLAS may want to be separated to low mu.
ALICE will request a static separation during the scans of ATLAS and CMS. The target is still to be defined.
The DCCT24bit will be used as the standard current monitor for the scan. This device does not require a setup of the range (since its dynamic range is large enough). However the DCCT should be used as a backup and a cross check (when the intensitites of the beams are small enough). Its range should be blocked to "range 3". Enrico Bravin remarked that the DCCT values of all ranges can always be found in Timber and that only the DIP value is blocked to one range.
Filling schemes for both fills will be prepared by Massimiliano Ferro-Luzzi.
A short discussion on wether or not to switch on the Orbit Feedback System (OFB) took place: In 2015 the OFB was switched on but this year the orbit stability is much better. Jörg Wenninger commented that the OFB should be switched off to be sure to have no interference during the scans. It might be switched on between scans. Witold remarked that you might not want to switch it on between x- and y-scans but prefer to have it on all time.